OpenStudy (anonymous):

Which expression is equivalent to the given expression? 13(x – 21) A. 13x – 21 B. 13x – 34 C. 13x – 13 • 21 D. 13 – 21x

jimthompson5910 (jim_thompson5910):

hints: a*(b+c) = a*b + a*c a*(b-c) = a*b - a*c

OpenStudy (anonymous):

c?

jimthompson5910 (jim_thompson5910):

good, you multiply that outer 13 by everything inside the parenthesis and subtract

OpenStudy (anonymous):

Which expression is equivalent to the given expression? 17x – 32x A. 17(x – 32) B. 17x(1 – 32) C. x(32 – 17) D. x(17 – 32)

OpenStudy (anonymous):

a?

jimthompson5910 (jim_thompson5910):

no, if you distributed in choice A, you'd get 17*x - 17*32

jimthompson5910 (jim_thompson5910):

but you want 17x - 32x

OpenStudy (anonymous):

d!

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

Identify the like terms. 7c, –7d, 7cd, –12cd A. 7c, 7cd B. 7cd, –12cd C. 7c, –7d D. 7c, –7d, 7cd

OpenStudy (anonymous):

b?

jimthompson5910 (jim_thompson5910):

yep, basically the same variable terms (both have cd) the others have a mix of letters but not the same mix

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

no d?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yeah you got it correct with b

OpenStudy (anonymous):

Combine like terms. 4s – 15s A. 11s B. –11s C. –19s D. 19s

OpenStudy (anonymous):

d?

jimthompson5910 (jim_thompson5910):

think of it like 4-15 then tack on 's' when you're done subtracting

OpenStudy (anonymous):

11 s?

OpenStudy (anonymous):

or would it be negitive?

jimthompson5910 (jim_thompson5910):

negative it's like you have 4 dollars and you buy a 15 dollar item, you'll have -11 dollars...ie be 11 dollars in debt

OpenStudy (anonymous):

Evaluate. 2 – 4 • 3 + 6 • 3 A. –28 B. 0 C. 8 D. 12

OpenStudy (anonymous):

-28?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you'll use PEMDAS here

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 3 – 7x + 2x when x = 3 A. –12 B. –18 C. –30 D. –24 <----

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

did you replace every x with 3, then evaluate?

jimthompson5910 (jim_thompson5910):

3 – 7x + 2x = 3 – 7*(3) + 2*(3) = ???

OpenStudy (anonymous):

18?

jimthompson5910 (jim_thompson5910):

3 – 7*(3) + 2*(3) = 3 - 21 + 6 = ??

jimthompson5910 (jim_thompson5910):

refresh the page if you get weird looking symbols showing up

OpenStudy (anonymous):

its 24

jimthompson5910 (jim_thompson5910):

no it's not

jimthompson5910 (jim_thompson5910):

hopefully you aren't just guessing

OpenStudy (anonymous):

im not i swear

jimthompson5910 (jim_thompson5910):

think of 3-21 as 21-3 just make the result negative

jimthompson5910 (jim_thompson5910):

so a similar example 2-10 = -8 notice how 10-2 = 8 and I just made the result negative

OpenStudy (anonymous):

-18 ?

jimthompson5910 (jim_thompson5910):

yeah 3 - 21 is -18

jimthompson5910 (jim_thompson5910):

3 - 21 + 6 = -18 + 6 = ??

jimthompson5910 (jim_thompson5910):

keep in mind that 18 is negative

OpenStudy (anonymous):

okay thanks

jimthompson5910 (jim_thompson5910):

so what is -18+6 equal to?

jimthompson5910 (jim_thompson5910):

the answer isn't -18 btw

OpenStudy (anonymous):

i guessed i dont care its 11 here my freinds in the hospital bad crash and i really dont care to do it sorry thanks tho bye